OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral surface area? Use π = 22/7

OpenStudy (lgbasallote):

lateral surface area = perimeter of base x height base area of cyclinder = pi (r^2) A/pi = r^2 r = sqrt (A/pi) r = sqrt (154/22/7) base perimeter = \[2 \pi rh\]=\[\large 2(19) (\frac{22}{7}) \sqrt{\frac {154}{\frac {22}{7}}}\] if we perform rationalization..we'll have \[\large 2(19) (\frac{22}{7}) \frac{\sqrt{(154)(\frac{22}{7})}}{\frac{22}{7}}\] so 22/7 gets caceled out...we're left with \[\large 2(19) \sqrt{(154)(\frac{22}{7})}\] i think that can be solved via calculator right? :D

OpenStudy (anonymous):

I cant do that with my calculator...